OpenStudy (anonymous):

Tennis balls are packed into a square box 12 in. on one side. The tennis balls can be touching, but they can’t be compressed. The remaining space is filled with packing beads. How many tennis balls can be packed into the box? What is the volume occupied by the packing beads? What percentage of the volume of the box is filled with beads?

OpenStudy (anonymous):

try this b) What is the velocity of each ball as it strikes the ground? ball thrown downward m/s ball thrown upward m/s (c) How far apart are the balls 0.800 s after they are thrown? * physics - tchrwill, Thursday, March 3, 2011 at 10:57am Thrown down: From Vf^2 = Vo^2 + 2gh, Vf = 24.89m/s. From Vf = Vo + gt, t = 1.04sec. Thrown upward: From Vf = Vo - gt, t = 1.5sec. to Vf = 0 From h = Vo(t) - gt^2/2, h = 11.025m. From peak height to ground = 31.625m. From Vf^2 = Vo^2 + 2gh, Vf = 24.896m/s. Distance traveled by ball thrown upward after .8 sec. derives from d = Vo(t) - gt^2/2 or d = +8.624m. Distance traveled by ball thrown downward after .8 sec. derives from d = Vo(t) + gt^2/2 or d = 14.89m. a) Difference in air time is 4.04 - 1.04 = 3sec. b) Vf thrown downward = 24.89m/s ...Vf thrown upward = 24.89m/s c) Therefore, after .8 sec., the balls are 23.52m apart.